jagr2713 (jagr2713):

Let f(x) = x2 - 8x + 5. Find f(-1). -4 -3 13 14

OpenStudy (ankitshaw):

u just need to put -1 in place of x and solve it

jagr2713 (jagr2713):

no its Let f(x) = x^2- 8x + 5. Find f(-1). -4 -3 13 14

OpenStudy (ankitshaw):

see f(x) = x^2 - 8x + 5 so f(-1) = (-1)^2 - 8(-1) +5 ... i hope now u understood

jagr2713 (jagr2713):

oh ok

jagr2713 (jagr2713):

so its D

OpenStudy (shubhamsrg):

yes it D ^_^

OpenStudy (shubhamsrg):

its*